# Critical Batch Size Revisited: A Simple Empirical Approach to Large-Batch Language Model Training

### v2

@@ Abstract
The right batch size is important when training language models at scale: a large batch size is necessary for fast training, but a batch size that is too large will harm token efficiency. To navigate this tradeoff, McCandlish et al. (2018) suggest that a critical batch size (CBS), below which training will not substantially degrade loss, can be estimated based on the gradient noise scale during training. While their method has been adopted in practice, e.g., when training GPT-3, strong assumptions are required to justify gradient noise as a proxy for the CBS, which makes it unclear whether their approach should be trusted in practice, limiting its applicability. In this paper, we introduce a simple, empirical approach to directly measure the CBS and show how the CBS evolves over training. Applying our approach to the OLMo models, we find that CBS is near 0 at initialization, increases rapidly at first, and then plateaus as training progresses. Furthermore, we find that this trend holds across different model sizes (1B and 7B), suggesting CBS from small training runs can inform larger-scale training runs. Our findings about how the CBS changes over training motivate batch size warmup as a natural way to reliably train language models at large batch size: start the batch size small and increase it as the CBS grows. To validate this claim, we use batch size warmup to train OLMo 1B to slightly better loss than the original training run with 43% fewer gradient steps. This shows how our framework can be applied to reliably train language models at larger batch sizes, increasing data parallelism without compromising performance.
@@ 1 Introduction
Critical Batch Size Hypothesis (McCandlish et al., 2018) There is some critical batch size $B^{*}$ up to which increasing the batch size (and appropriately modifying the learning rate) approximately preserves the loss trajectory as a function of tokens trained, but, above which, the loss trajectory degrades.
If such a CBS $B^{*}$ exists (and we can measure it), it represents a reasonable balance between efficiency and performance (i.e., loss) and is thus a practically useful batch size at which to train. Working in a simplified theoretical setup, McCandlish et al. (2018) derive a correspondence between the CBS and the gradient noise scale, i.e., the variance of the per-example gradients from the training distribution. They suggest that an estimator for the gradient noise scale should be used in practice as a proxy for the CBS, and this can in turn be used to set the batch size for large-scale pretraining runs. The noise scale also appears to have been adopted in practice as a proxy for the CBS, having been mentioned explicitly in the GPT-3 technical report (Brown et al., 2020), and inspiring a flurry of methodological innovations for better noise scale estimation (Gray et al., 2023, 2024).
While appealing, the link between the gradient noise and the CBS requires several strong assumptions to justify: specifically, it assumes the SGD optimizer and that gradients are well-conditioned (cf. Section 2). Thus, it is unclear whether the noise scale should be a meaningful proxy for the CBS for language model pretraining in practice, where the Adam optimizer is often used and the optimization may not be well-conditioned. To this end, we aim to address the following practical questions that remain for effectively leveraging the CBS viewpoint to train language models at larger batch sizes:
@@
This work documents our attempts to answer these questions in order to operationalize the CBS for large-batch training. We focus our investigation on the OLMo models (Groeneveld et al., 2024; OLMo et al., 2025), due to their open weights and data, making the following contributions:
@@
    We use our method to study how our CBS measurement changes over the course of training, finding it improves rapidly initially but than flattens off. Further, we find that the CBS not depend strongly on model size, in line with past findings using different methodology (Zhang et al., 2024).
@@ 2 Background: Estimating CBS via Gradient Noise Scale
Past empirical work has aimed to measure the CBS by launching many training runs to the same target loss, which is expensive (Zhang et al., 2019, 2024), or by using the gradient noise scale as a proxy (McCandlish et al., 2018; Gray et al., 2023, 2024). In contrast, we will introduce a new method that uses a small amount of additional training to estimate the CBS, which is less expensive than launching many full training runs and does not make any strong assumptions like the noise scale. Before introducing our method, we review the noise scale framework used to estimate the CBS in prior work and the underlying assumptions it relies on.
@@
McCandlish et al. (2018) suggest that the CBS can be measured in terms of the gradient noise scale, i.e., the variance of the gradients within a batch. Concretely, their recommendations to measure the CBS and adapt the learning rate are as follows:
@@
Existing Method: Noise Scale Proxy for CBS (McCandlish et al., 2018) Let $G$ be the full gradient and let $\Sigma$ be the covariance matrix for the gradient across data examples. We first compute $\mathcal{B}_{\mathsf{simple}}$ as a proxy for the CBS (using an efficient statistical estimator): $\mathcal{B}_{\mathsf{simple}}=\frac{\mathrm{tr}(\Sigma)}{\lVert G\rVert^{2}}.$ We set $B_{\mathsf{simple}}$ as modified batch size $B^{*}$ and linearly scale the learning rate $\eta^{*}$: $B^{*}$ $=\mathcal{B}_{\mathsf{simple}}$ (1) $\eta^{*}$ $=\frac{B^{*}}{B}\cdot\eta.$ (2)
@@
This viewpoint is attractive due to its simplicity and tractability, and it has inspired improved methods for estimating the noise scale (Gray et al., 2023, 2024). However, the link between noise scale and CBS requires several strong assumptions to justify, which we will argue motivates revisiting other approaches for measuring the CBS. McCandlish et al. (2018) consider training a model on a loss surface where the loss landscape is well approximated by its second-order Taylor expansion. The first crucial assumption in their analysis is that optimizer used to decrease the loss is SGD:
@@
Assumption 1: SGD Optimizer (McCandlish et al., 2018) The step taken to reduce the loss is a noisy estimate of the true gradient, computed via $B$ samples.
@@
This assumption may seem benign, but it is worth noting that it is not typically met in practice, as LMs are trained with the Adam optimizer (Kingma and Ba, 2017). Moreover, by analyzing training dynamics in terms of stochastic differential equations (Li et al., 2021), Malladi et al. (2022) suggest that Equation 2 is an appropriate scaling rule for SGD, but not for Adam, where a quadratic scaling rule is more principled. Thus, if training with Adam, it is unclear whether the noise scale should be a valid proxy for CBS, and, if so, how to operationalize the scaling rule for learning rate.
@@
McCandlish et al. (2018) also require another strong assumption to derive the noise scale method for estimating the CBS. In general, their noise-scale-based estimate of the CBS has a more complex form than $\mathcal{B}_{\mathrm{simple}}$ involving the Hessian $H$:
@@
In order to justify that $B^{*}$ can be computed as $\mathcal{B}_{\mathrm{simple}}$, McCandlish et al. (2018) assume:
Assumption 2: Well-Conditioned Optimization (McCandlish et al., 2018) The Hessian $H$ is a multiple of the identity matrix. It follows that $\mathcal{B}_{\mathrm{noise}}=\mathcal{B}_{\mathrm{simple}}$.
This is a strong assumption required to justify using $\mathcal{B}_{\mathrm{simple}}$ as a proxy for the CBS because computing Hessians would be too expensive to be practice. McCandlish et al. (2018) suggest informally that, without this assumption, $\mathcal{B}_{\mathrm{simple}}$ may still be correlated with $\mathcal{B}_{\mathrm{noise}}$, but it is not obvious why this should be the case. Even if this is true, it still poses a real problem for the noise scale methodology, since the goal of the method is to produce an absolute measure of $B^{*}$. It is unclear for practitioners what coefficient should be used to translate $\mathcal{B}_{\mathrm{simple}}$ to $B^{*}$—and, more fundamentally, whether it is even valid to assume that such a coefficient exists.
@@ 3 Our Method: Measuring the CBS via Local Branched Training
As discussed in Section 2, using the gradient noise scale $\mathcal{B}_{\textrm{simple}}$ as a proxy to estimate the CBS relies on several strong assumptions. In light of this, it unclear whether we can trust the gradient noise scale as a proxy for CBS. We thus argue that we should instead aim to measure the CBS directly, without the need for any strong assumptions to justify an indirect proxy. After introducing our measurement approach in this section, we will show in Section 4 how these measurements can be applied to train language models to the same (or better) target loss with fewer gradients steps.
@@ 3.1 Method
We introduce a simple branched training method that directly approximates the CBS by launched branched training runs from a checkpoint, which allows us to identify $B^{*}$ as the largest batch size that does not degrade in loss relative to smaller batch sizes as visualized in Figure 2. To make this tractable, we train only for a fixed token budget $\Delta$, assuming that if $B^{*}$ recovers in loss by $\Delta$, its loss will continue to match smaller batch sizes onwards as well. This allows us to estimate the CBS with only a small amount of additional training (controlled by $\Delta$). Further, as we will find later in Figure 2, the CBS trend remains consistent across model sizes, so CBS measurements with small models could be used to inform large-scale training runs.
@@
Figure 1: Smoothed final loss after branched training at particular checkpoints, with $B^{*}$ shown as the dotted red line. Each point represents the loss achieved by a specific branched training run after 2B tokens. Our method detects the point at which loss starts to increase, heuristically tolerating noise within $\epsilon=0.01$. These plots show how this plays out for three particular checkpoints; see Appendix A for loss curves for all checkpoints.
@@
    Loss Tolerance. Viewing loss as a function of batch size multiplier $k$ (cf.Figure 1), we need a way to determine whether the loss at $k^{*}$ has increased relative to all $k<k^{*}$. We operationalize this with a tolerance parameter $\epsilon$, which we set to 0.01 arbitrarily. In principle, tolerance could be set in a more principled way using a statistical test in future work.
@@
It is worth comparing our method to other work that empirically measures the CBS. Our method of training for a fixed token budget $\Delta$ and measuring the change in loss can be understood as the dual view to measuring the number of steps required to achieve a target loss, which has been used in prior work (Zhang et al., 2019, 2024). Reformulating the measurement in this way has the nice property that the training budget can be fixed in advance. Under the local recovery assumption, it also allows us to train for only $\Delta$ tokens, which means we do not have to launch full training runs for each batch size. Finally, unlike Zhang et al. (2024), we apply our method at various checkpoints throughout training, whereas they apply it only from initialization. This means we can estimate the local CBS at a specific point in training rather than just the global CBS.
@@ 3.2 Experimental Setup
We aim to measure the CBS over the course of model training and the role of model size. Because our method requires pretraining checkpoints and access to the pretraining data, we use the OLMo 1B and OLMo 7B models for our experiments (OLMo et al., 2025), whose pretraining data is openly available (Soldaini et al., 2024). For each model, we take a variety of checkpoints over the course of training, allowing us to assess how the CBS changes over the course of training; see Appendix A for more details. We also compute the noise scale across training checkpoints using the estimator proposed by McCandlish et al. (2018) to assess whether it is a valid proxy for the CBS we measure.
We define the interval for the CBS at each checkpoint by choosing $B^{*}$ (as defined in Section 3) as a lower bound, and by picking the least $k>k^{*}$ as an upper bound. The plotted point represents the geometric mean of these interval endpoints. We measure the CBS in documents, with a pretraining sequence length of 4096 tokens per document.
@@ 3.3 Results: CBS Over Training and Across Model Sizes
Figure 2 shows the CBS $B^{*}$ measured via Section 3.1 across training checkpoints for OLMo 1B and 7B. The CBS increases over training in a similar way for both model sizes: the CBS starts near 0, grows rapidly within the first 50k tokens, and then plateaus around 4096.
@@
Prior work has suggested that the CBS is largely independent of model size, scaling primarily with data size (Zhang et al., 2024; Bergsma et al., 2025). This is largely consistent with our findings, as the CBS curves like qualitatively similar at the 1B and 7B scales. In addition, the fact that the CBS grows over training suggests that the “aggregate” CBS should also increase as we train on more data since the average CBS over the course of training should increase. We elaborate on this in Appendix D: while CBS growth over training predicts that aggregate CBS should increase with data size, it is unclear whether the CBS growth pattern we observe would predict the aggregate CBS $\propto\sqrt{D}$ scaling law found in prior work (Zhang et al., 2024; Bergsma et al., 2025).
@@
As discussed in Section 2, the gradient noise scale has been proposed as a proxy to measure the CBS (McCandlish et al., 2018), though this connection relies on strong assumptions to justify. We thus empirically compare our measurement of the CBS to the gradient noise scale. Figure 3 shows that, for both OLMo 1B and 7B, the gradient noise scale underestimates the CBS by several orders of magnitude. Furthermore, especially for OLMo 7B, the qualitative trend does not match the CBS. For OLMo 1B, the qualitative pattern is more similar. However, since this similarity is not found for both models, we conclude that, in general, the noise scale cannot be used reliably as a proxy for the CBS.
Figure 3: Gradient noise scale for OLMo 1B and 7B computed via the estimator of McCandlish et al. (2018) with 95% confidence intervals; details in Appendix B. The gradient noise scale underestimates the CBS (cf. Figure 2) and the qualitative trend does not clearly match, especially for OLMo 7B.

### v3

@@ Abstract
The right batch size is important when training language models at scale: a large batch size is necessary for fast training, but a batch size that is too large will harm token efficiency. To navigate this tradeoff, mccandlish2018empirical suggest that a critical batch size (CBS), below which training will not substantially degrade loss, can be estimated based on the gradient noise scale during training. While their method has been adopted in practice, e.g., when training GPT-3, strong assumptions are required to justify gradient noise as a proxy for the CBS, which makes it unclear whether their approach should be trusted in practice, limiting its applicability. In this paper, we introduce a simple, empirical approach to directly measure the CBS and show how the CBS evolves over training. Applying our approach to the OLMo models, we find that CBS is near 0 at initialization, increases rapidly at first, and then plateaus as training progresses. Furthermore, we find that this trend holds across different model sizes (1B and 7B), suggesting CBS from small training runs can inform larger-scale training runs. Our findings about how the CBS changes over training motivate batch size warmup as a natural way to reliably train language models at large batch size: start the batch size small and increase it as the CBS grows. To validate this claim, we use batch size warmup to train OLMo 1B to slightly better loss than the original training run with 43% fewer gradient steps. This shows how our framework can be applied to reliably train language models at larger batch sizes, increasing data parallelism without compromising performance.
@@ 1 Introduction
Critical Batch Size Hypothesis (mccandlish2018empirical) There is some critical batch size $B^{*}$ up to which increasing the batch size (and appropriately modifying the learning rate) approximately preserves the loss trajectory as a function of tokens trained, but, above which, the loss trajectory degrades.
If such a CBS $B^{*}$ exists (and we can measure it), it represents a reasonable balance between efficiency and performance (i.e., loss) and is thus a practically useful batch size at which to train. Working in a simplified theoretical setup, mccandlish2018empirical derive a correspondence between the CBS and the gradient noise scale, i.e., the variance of the per-example gradients from the training distribution. They suggest that an estimator for the gradient noise scale should be used in practice as a proxy for the CBS, and this can in turn be used to set the batch size for large-scale pretraining runs. The noise scale also appears to have been adopted in practice as a proxy for the CBS, having been mentioned explicitly in the GPT-3 technical report (brown2020language), and inspiring a flurry of methodological innovations for better noise scale estimation (gray2023efficient; gray2024normalization).
While appealing, the link between the gradient noise and the CBS requires several strong assumptions to justify: specifically, it assumes the SGD optimizer and that gradients are well-conditioned (cf. Section˜2). Thus, it is unclear whether the noise scale should be a meaningful proxy for the CBS for language model pretraining in practice, where the Adam optimizer is often used and the optimization may not be well-conditioned. To this end, we aim to address the following practical questions that remain for effectively leveraging the CBS viewpoint to train language models at larger batch sizes:
@@
This work documents our attempts to answer these questions in order to operationalize the CBS for large-batch training. We focus our investigation on the OLMo models (groeneveld-etal-2024-olmo; olmo20252olmo2furious), due to their open weights and data, making the following contributions:
@@
    We use our method to study how our CBS measurement changes over the course of training, finding it improves rapidly initially but than flattens off. Further, we find that the CBS not depend strongly on model size, in line with past findings using different methodology (zhang2024cbs).
@@ 2 Background: Estimating CBS via Gradient Noise Scale
Past empirical work has aimed to measure the CBS by launching many training runs to the same target loss, which is expensive (zhang-2019-algorithmic; zhang2024cbs), or by using the gradient noise scale as a proxy (mccandlish2018empirical; gray2023efficient; gray2024normalization). In contrast, we will introduce a new method that uses a small amount of additional training to estimate the CBS, which is less expensive than launching many full training runs and does not make any strong assumptions like the noise scale. Before introducing our method, we review the noise scale framework used to estimate the CBS in prior work and the underlying assumptions it relies on.
@@
mccandlish2018empirical suggest that the CBS can be measured in terms of the gradient noise scale, i.e., the variance of the gradients within a batch. Concretely, their recommendations to measure the CBS and adapt the learning rate are as follows:
@@
Existing Method: Noise Scale Proxy for CBS (mccandlish2018empirical) Let $G$ be the full gradient and let $\Sigma$ be the covariance matrix for the gradient across data examples. We first compute $\mathcal{B}_{\mathsf{simple}}$ as a proxy for the CBS (using an efficient statistical estimator): $\mathcal{B}_{\mathsf{simple}}=\frac{\mathrm{tr}(\Sigma)}{\lVert G\rVert^{2}}.$ We set the modified batch size to $\mathcal{B}_{\mathsf{simple}}$ and linearly scale the learning rate $\eta^{*}$: $B^{*}$ $=\mathcal{B}_{\mathsf{simple}}$ (1) $\eta^{*}$ $=\frac{B^{*}}{B}\cdot\eta.$ (2)
@@
This viewpoint is attractive due to its simplicity and tractability, and it has inspired improved methods for estimating the noise scale (gray2023efficient; gray2024normalization). However, the link between noise scale and CBS requires several strong assumptions to justify, which we will argue motivates revisiting other approaches for measuring the CBS. mccandlish2018empirical consider training a model on a loss surface where the loss landscape is well approximated by its second-order Taylor expansion. The first crucial assumption in their analysis is that optimizer used to decrease the loss is SGD:
@@
Assumption 1: SGD Optimizer (mccandlish2018empirical) The step taken to reduce the loss is a noisy estimate of the true gradient, computed via $B$ samples.
@@
This assumption may seem benign, but it is worth noting that it is not typically met in practice, as LMs are trained with the Adam optimizer (kingma-2017-adam). Moreover, by analyzing training dynamics in terms of stochastic differential equations (li2021validity), malladi2022sdes argue theoretically that Equation˜2 is an appropriate scaling rule for SGD, but not for Adam, where a square-root scaling rule is more principled. The square-root scaling for Adam is also supported by the theoretical analysis of li2024surge. Thus, when training with Adam, it seems that the linear scaling rule assumed by mccandlish2018empirical should not apply.
@@
A more fundamental issue is that mccandlish2018empirical also require another strong assumption to derive the noise scale method for estimating the CBS. In general, their noise-scale-based estimate of the CBS has a more complex form than $\mathcal{B}_{\mathrm{simple}}$ involving the Hessian $H$:
@@
In order to justify that $B^{*}$ can be computed as $\mathcal{B}_{\mathrm{simple}}$, mccandlish2018empirical assume:
Assumption 2: Well-Conditioned Optimization (mccandlish2018empirical) The Hessian $H$ is a multiple of the identity matrix. It follows that $\mathcal{B}_{\mathrm{noise}}=\mathcal{B}_{\mathrm{simple}}$.
This is a strong assumption required to justify using $\mathcal{B}_{\mathrm{simple}}$ as a proxy for the CBS because computing Hessians would be too expensive to be practice. mccandlish2018empirical suggest informally that, without this assumption, $\mathcal{B}_{\mathrm{simple}}$ may still be correlated with $\mathcal{B}_{\mathrm{noise}}$, but it is not obvious why this should be the case. Even if this is true, it still poses a real problem for the noise scale methodology, since the goal of the method is to produce an absolute measure of $B^{*}$. It is unclear for practitioners what coefficient should be used to translate $\mathcal{B}_{\mathrm{simple}}$ to $B^{*}$—and, more fundamentally, whether it is even valid to assume that such a coefficient exists.
@@ 3 Our Method: Measuring the CBS via Local Branched Training
As discussed in Section˜2, using the gradient noise scale $\mathcal{B}_{\textrm{simple}}$ as a proxy to estimate the CBS relies on several strong assumptions. In light of this, it unclear whether we can trust the gradient noise scale as a proxy for CBS. We thus argue that we should instead aim to measure the CBS directly, without the need for any strong assumptions to justify an indirect proxy. After introducing our measurement approach in this section, we will show in Section˜4 how these measurements can be applied to train language models to the same (or better) target loss with fewer gradients steps.
@@ 3.1 Method
We introduce a simple branched training method that directly approximates the CBS by launched branched training runs from a checkpoint, which allows us to identify $B^{*}$ as the largest batch size that does not degrade in loss relative to smaller batch sizes as visualized in Figure˜2. To make this tractable, we train only for a fixed token budget $\Delta$, assuming that if $B^{*}$ recovers in loss by $\Delta$, its loss will continue to match smaller batch sizes onwards as well. This allows us to estimate the CBS with only a small amount of additional training (controlled by $\Delta$). Further, as we will find later in Figure˜2, the CBS trend remains consistent across model sizes, so CBS measurements with small models could be used to inform large-scale training runs.
@@
Figure 1: Smoothed final loss after branched training at particular checkpoints, with $B^{*}$ shown as the dotted red line. Each point represents the loss achieved by a specific branched training run after 2B tokens. Our method detects the point at which loss starts to increase, heuristically tolerating noise within $\epsilon=0.01$. These plots show how this plays out for three particular checkpoints; see Appendix˜A for loss curves for all checkpoints.
@@
    Loss Tolerance. Viewing loss as a function of batch size multiplier $k$ (cf.Figure˜1), we need a way to determine whether the loss at $k^{*}$ has increased relative to all $k<k^{*}$. We operationalize this with a tolerance parameter $\epsilon$, which we set to 0.01 arbitrarily. In principle, tolerance could be set in a more principled way using a statistical test in future work.
@@
It is worth comparing our method to other work that empirically measures the CBS. Our method of training for a fixed token budget $\Delta$ and measuring the change in loss can be understood as the dual view to measuring the number of steps required to achieve a target loss, which has been used in prior work (zhang-2019-algorithmic; zhang2024cbs). Reformulating the measurement in this way has the nice property that the training budget can be fixed in advance. Under the local recovery assumption, it also allows us to train for only $\Delta$ tokens, which means we do not have to launch full training runs for each batch size. Finally, unlike zhang2024cbs, we apply our method at various checkpoints throughout training, whereas they apply it only from initialization. This means we can estimate the local CBS at a specific point in training rather than just the global CBS.
@@ 3.2 Experimental Setup
We aim to measure the CBS over the course of model training and the role of model size. Because our method requires pretraining checkpoints and access to the pretraining data, we use the OLMo 1B and OLMo 7B models for our experiments (olmo20252olmo2furious), whose pretraining data is openly available (soldaini2024dolma). For each model, we take a variety of checkpoints over the course of training, allowing us to assess how the CBS changes over the course of training; see Appendix˜A for more details. We also compute the noise scale across training checkpoints using the estimator proposed by mccandlish2018empirical to assess whether it is a valid proxy for the CBS we measure.
We define the interval for the CBS at each checkpoint by choosing $B^{*}$ (as defined in Section˜3) as a lower bound, and by picking the least $k>k^{*}$ as an upper bound. The plotted point represents the geometric mean of these interval endpoints. We measure the CBS in documents, with a pretraining sequence length of 4096 tokens per document.
@@ 3.3 Results: CBS Over Training and Across Model Sizes
Figure˜2 shows the CBS $B^{*}$ measured via Section˜3.1 across training checkpoints for OLMo 1B and 7B. The CBS increases over training in a similar way for both model sizes: the CBS starts near 0, grows rapidly within the first 50k tokens, and then plateaus around 4096.
@@
Prior work has suggested that the CBS is largely independent of model size, scaling primarily with data size (zhang2024cbs; bergsma-2025-powerlines). This is largely consistent with our findings, as the CBS curves like qualitatively similar at the 1B and 7B scales. In addition, the fact that the CBS grows over training suggests that the “aggregate” CBS should also increase as we train on more data since the average CBS over the course of training should increase. We elaborate on this in Appendix˜D: while CBS growth over training predicts that aggregate CBS should increase with data size, it is unclear whether the CBS growth pattern we observe would predict the aggregate CBS $\propto\sqrt{D}$ scaling law found in prior work (zhang2024cbs; bergsma-2025-powerlines).
@@
As discussed in Section˜2, the gradient noise scale has been proposed as a proxy to measure the CBS (mccandlish2018empirical), though this connection relies on strong assumptions to justify. We thus empirically compare our measurement of the CBS to the gradient noise scale. Figure˜3 shows that, for both OLMo 1B and 7B, the gradient noise scale underestimates the CBS by several orders of magnitude. Furthermore, especially for OLMo 7B, the qualitative trend does not match the CBS. For OLMo 1B, the qualitative pattern is more similar. However, since this similarity is not found for both models, we conclude that, in general, the noise scale cannot be used reliably as a proxy for the CBS.
Figure 3: Gradient noise scale for OLMo 1B and 7B computed via the estimator of mccandlish2018empirical with 95% confidence intervals; details in Appendix˜B. The gradient noise scale underestimates the CBS (cf. Figure˜2) and the qualitative trend does not clearly match, especially for OLMo 7B.